OpenStudy (anonymous):

solve the equation \[\frac{ x }{ x+2 } = \frac{ 9 }{ x+9 }\]

OpenStudy (anonymous):

Solve for X

OpenStudy (kainui):

Where are you getting stuck?

OpenStudy (anonymous):

I wasnt sure what it was really asking me, so solve for x?

OpenStudy (anonymous):

Consider anything added by X as a single entity so as you may notice by examining the denominators you have 2 different fractions. Make them similar and solve for X

OpenStudy (anonymous):

Hint: Common denominators

OpenStudy (anonymous):

this would be easier with cross multiplication

OpenStudy (anonymous):

these are the answers in the mutiple choice, im kind of confused by them \[A) X= \frac{ -9\pm2\sqrt{9} }{ 2 }\] B) x=2, x= 9 C) no solution D) \[\pm3\sqrt{2}\]

OpenStudy (anonymous):

i got it, thanks guys

OpenStudy (anonymous):

its D?